OpenStudy (anonymous):

Solve for x. Don't forget to perform a check. 5/11x=20

OpenStudy (anonymous):

x = 1/44

OpenStudy (anonymous):

no direct answers...pleez

OpenStudy (anonymous):

can you please give me the steps

OpenStudy (anonymous):

5/11x = 20 multiply both sides by 11x 5 = 11x * 20 = 220x divide both sides by 220 x = 5/220 = 1/44